Solve the polyomial equation.
X^3-6x^2+6x=0
Show step by step
Thanks in advance.

Answer by jim_thompson5910(35256)

Hint:

x^3 - 6x^2 + 6x = 0

x(x^2 - 6x + 6) = 0

x = 0 or x^2 - 6x + 6 = 0

So one solution or root is x = 0

The other two roots can be found if you solve x^2 - 6x + 6 = 0

Use the quadratic formula to solve x^2 - 6x + 6 = 0 for x